# Kaon-induced production of strange hidden-charm molecular pentaquarks $P^{\Lambda}_{\psi s}$ from proton

### v1

@@ III.1 Molecular states without coupled-channel effects
Five poles are identified within the energy range of 4200 to 4550 MeV based on the interactions considered. Since the $\bar{K}N$ threshold is significantly below the charmed energy region and the couplings between channels are turned off, the results closely resemble those from our previous work where the $\bar{K}N$ channel was excluded Zhu:2022wpi . These poles correspond to five molecular states generated from the single-channel interactions of $\Xi_{c}^{*}\bar{D}$ with spin parity $J^{P}=3/2^{-}$, $\Xi_{c}\bar{D}^{*}$ with $1/2^{-}$ and $3/2^{-}$, $\Xi^{\prime}_{c}\bar{D}$ with $1/2^{-}$, and $\Xi_{c}\bar{D}$ with $1/2^{-}$. As the cutoff increases, all poles shift further away from their respective thresholds. Two states with spin parities $1/2^{-}$ and $3/2^{-}$ from the $\Xi_{c}\bar{D}^{*}$ interaction have nearly the same mass as those found in the calculation without coupled-channel effects, approximately 4474 MeV, which is slightly higher than the mass of $P^{\Lambda}_{\psi s}(4459)$. The $\Xi_{c}\bar{D}(1/2^{-})$ interaction generates an S-wave bound state with a mass around 4335 MeV, closely matching the mass of $P^{\Lambda}_{\psi s}(4338)$. The $\Xi^{\prime}_{c}\bar{D}$ interaction produces a molecular state with spin parity $1/2^{-}$. The $\Xi_{c}^{*}\bar{D}$ channel gives rise to a molecular state with spin parity $3/2^{-}$. Even when the cutoff parameter is adjusted to its maximum within a reasonable range, no bound states are formed for the channels $\Lambda_{c}\bar{D}_{s}^{*}$ and $\Lambda_{c}\bar{D}_{s}$.
III.2 Cross section for $J^{P}=1/2^{-}$ in energy region $[4350-4550]$ MeV
@@
First, we consider the cross section for $J^{P}=1/2^{-}$ in the energy range $[4350-4550]$ MeV. In the calculation without coupled-channel effects, two poles with spin parity $1/2^{-}$ are generated in this energy region, $\Xi^{\prime}_{c}\bar{D}(1/2^{-})$ and $\Xi_{c}\bar{D}^{*}(1/2^{-})$, with the latter corresponding to $P^{\Lambda}_{\psi s}(4459)$. In Fig. 2, after including the coupled-channel effects, the poles for these two states are shown in the complex energy plane. The corresponding cross sections in the channels considered are also presented.
@@
The poles in the complex energy plane manifest as structures in the cross section, which can be calculated from the scattering amplitudes along the real axis. As shown in Fig. 2, the cross sections exhibit complex structures due to the coupled-channel interactions in the considered energy region. Generally, the cross sections for the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels are on the order of tens to hundreds of nanobarns (nb), significantly larger than those for the $\Xi_{c}\bar{D}$ and $\Xi^{\prime}_{c}\bar{D}$ channels, which are on the order of several nb due to different phase spaces and exchange particles as shown in Table 2. The cross section for the $\Lambda J/\psi$ channel is much smaller than the others, as it cannot couple directly with the $\bar{K}N$ channel.
@@
III.3 Cross section for $J^{P}=3/2^{-}$ in energy region $[4350-4550]$ MeV
@@
III.4 Cross section for $J^{P}=1/2^{-}$ in energy region $[4250-4380]$ MeV
The calculation without coupling effects suggests the presence of a bound state originating from the $\Xi_{c}\bar{D}$ interaction, corresponding to $P^{\Lambda}_{\psi s}(4338)$. As shown in Fig. 4, after accounting for coupled-channel effects, the state acquires a small width. When the cutoff parameter $\Lambda$ increases, the pole shifts leftward along the real axis, and the pole near the $\Xi_{c}\bar{D}$ threshold becomes more prominent.
@@ IV Summary
These findings suggest that the strange hidden-charmed pentaquarks, such as $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$, may potentially be observed in kaon-induced production processes. High-precision experiments utilizing kaon beams would be instrumental in confirming the existence of these states and exploring their properties in detail. Such studies could provide valuable insights into the formation mechanisms of strange hidden-charm pentaquark.
V Acknowledgement
J. He is supported by the National Science Foundation of China (Grant No. 12475080), and J. T. Zhu is supported by the Start-up Funds of Changzhou University (Grant No. ZMF24020043) and the National Science Foundation of China (Grant No. 12405090).

### v2

@@ 3.1 Molecular states without coupled-channel effects
Five poles are identified within the energy range of 4200 to 4550 MeV based on the interactions considered. All poles identified in this work are located below threshold, meaning they reside on the Riemann sheet corresponding to the default propagator $G(z)$ of the respective channel. Since the $\bar{K}N$ threshold is significantly below the charmed energy region and the couplings between channels are turned off, the results closely resemble those from our previous work where the $\bar{K}N$ channel was excluded Zhu:2022wpi . These poles correspond to five molecular states generated from the single-channel interactions of $\Xi_{c}^{*}\bar{D}$ with spin parity $J^{P}=3/2^{-}$, $\Xi_{c}\bar{D}^{*}$ with $1/2^{-}$ and $3/2^{-}$, $\Xi^{\prime}_{c}\bar{D}$ with $1/2^{-}$, and $\Xi_{c}\bar{D}$ with $1/2^{-}$. As the cutoff increases, all poles shift further away from their respective thresholds. Two states with spin parities $1/2^{-}$ and $3/2^{-}$ from the $\Xi_{c}\bar{D}^{*}$ interaction have nearly the same mass as those found in the calculation without coupled-channel effects, approximately 4474 MeV, which is slightly higher than the mass of $P^{\Lambda}_{\psi s}(4459)$. The $\Xi_{c}\bar{D}(1/2^{-})$ interaction generates an S-wave bound state with a mass around 4335 MeV, closely matching the mass of $P^{\Lambda}_{\psi s}(4338)$. The $\Xi^{\prime}_{c}\bar{D}$ interaction produces a molecular state with spin parity $1/2^{-}$. The $\Xi_{c}^{*}\bar{D}$ channel gives rise to a molecular state with spin parity $3/2^{-}$. Even when the cutoff parameter is adjusted to its maximum within a reasonable range, no bound states are formed for the channels $\Lambda_{c}\bar{D}_{s}^{*}$ and $\Lambda_{c}\bar{D}_{s}$.
3.2 Cross section for $J^{P}=1/2^{-}$ in energy region $[4350-4550]$ MeV
@@
First, we consider the cross section for $J^{P}=1/2^{-}$ in the energy range $[4350-4550]$ MeV. In the calculation without coupled-channel effects, two poles with spin parity $1/2^{-}$ are generated in this energy region, $\Xi^{\prime}_{c}\bar{D}(1/2^{-})$ and $\Xi_{c}\bar{D}^{*}(1/2^{-})$, with the latter corresponding to $P^{\Lambda}_{\psi s}(4459)$. The imaginary part of the pole with a cutoff of 1.14 GeV is about 10 MeV, corresponding to a width of approximately 20 MeV, which is close to the value reported by LHCb, 17 MeV Aaij:2020gdg . In Fig. 2, after including the coupled-channel effects, the poles for these two states are shown in the complex energy plane. The corresponding cross sections in the channels considered are also presented.
@@
The poles in the complex energy plane manifest as structures in the cross section, which can be calculated from the scattering amplitudes along the real axis. As shown in Fig. 2, the cross sections exhibit complex structures due to the coupled-channel interactions in the considered energy region. Generally, the cross sections for the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels are on the order of tens to hundreds of nanobarns (nb), significantly larger than those for the $\Xi_{c}\bar{D}$ and $\Xi^{\prime}_{c}\bar{D}$ channels, which are on the order of several nb due to different phase spaces and exchange particles as shown in Table 2. The cross section for the $\Lambda J/\psi$ channel is on the order of 1nb or less, which is consistent with predictions from other theoretical studiesClymton:2021thh ; Paryev:2023icm ; Paryev:2022zdx ; Clymton:2022qlr . However, it is significantly smaller than those of other channels, as the $\Lambda J/\psi$ final state cannot couple directly to the $\bar{K}N$ channel.
@@
3.3 Cross section for $J^{P}=3/2^{-}$ in energy region $[4350-4550]$ MeV
@@
3.4 Cross section for $J^{P}=1/2^{-}$ in energy region $[4250-4380]$ MeV
The calculation without coupling effects suggests the presence of a bound state originating from the $\Xi_{c}\bar{D}$ interaction, corresponding to $P^{\Lambda}_{\psi s}(4338)$. As shown in Fig.4, after including coupled-channel effects, the state acquires a small imaginary part corresponding to a width of about 5 MeV, consistent with the LHCb observation LHCb:2022ogu . When the cutoff parameter $\Lambda$ increases, the pole shifts leftward along the real axis, and the pole near the $\Xi_{c}\bar{D}$ threshold becomes more prominent.
@@ 4 Summary
These findings suggest that the strange hidden-charmed pentaquarks, such as $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$, may potentially be observed in kaon-induced production processes. There are still many alternative assignments for the $P^{\Lambda}_{\psi s}$ Wang:2021itn ; Peng:2020hql ; Feijoo:2022rxf . Studying hidden-charmed pentaquarks through new production mechanisms, from both theoretical and experimental perspectives, may also be beneficial for distinguishing these different assignments. High-precision experiments utilizing kaon beams at J-PARC and JLab would be instrumental in confirming the existence of these states and exploring their properties in detail.
Acknowledgement J. He is supported by the National Science Foundation of China (Grant No. 12475080), and J. T. Zhu is supported by the Start-up Funds of Changzhou University (Grant No. ZMF24020043) and the National Science Foundation of China (Grant No. 12405090).
Data Availability Statement This manuscript has no associated data or the data will not be deposited. [Authors’ comment: This is a theoretical study and no external data are associated with this work.]